What does it mean to refer to a 20-in TV set or a 25-in TV set? Such units refer to the diagonal of the screen. A 10-in TV set also has a width of 8-in. What is its height?
Answer by checkley77(12844)

8^+W^2=10^
64=W^2=100
H^2=100-64
H^2=36
H=SQRT36
H=6 ANS. FOR THE HEIGHT.